OpenStudy (anonymous):

Solve algebraically: log15 + log(2x-3) - log2= 1/6

OpenStudy (anonymous):

use log properties log15+log(2x-3)=log(30x-45)

OpenStudy (anonymous):

You can then do log((30x-45)/2) = log(15x-45/2)=1/6 Raise both sides to the tenth to get rid of the log 15x-45/2=10^(1/6)

OpenStudy (anonymous):

thank you!